# AC conductivity and magnetic dichroism of two-dimensional antiferromagnetic Dirac semimetals

### v1

@@ I Introduction
On-demand manipulation and readout of the staggered Néel order parameter in antiferromagnetic (AFM) systems are challenging problems in the emerging field of AFM-based spintronics [1, 2]. Both theoretically and experimentally, spin-orbit and spin-transfer torques have been demonstrated to be finite in AFM systems under certain symmetry conditions [3, 4, 5, 6, 7, 8, 9, 10, 11]. It was shown that in AFM systems with preserved combined $\mathcal{PT}$ symmetry, while the time-reversal $\mathcal{T}$ symmetry and the inversion $\mathcal{P}$ symmetry are each broken, staggered spin-orbit torques acting on the magnetic sublattices may reorient the AFM Néel vector [11, 6, 5, 10, 3].
@@
Two-dimensional (2D) Dirac materials are a recently discovered class of 2D materials that are not gapped by spin-orbit couplings (SOCs) [12, 13, 14], and thus exhibit behavior distinct from graphene [15, 16], surface states of topological insulators [17], and 3D Dirac and Weyl semimetals [18, 19]. They have more than one Dirac point in their first Brillouin zone (BZ). Band crossing in Dirac points of these materials is protected by nonsymmorphic lattice symmetry [12]. The 2D CuMnAs is a prototype of Dirac materials with a long-range AFM order [20, 21, 22]. The recent discovery of electrical switching, via SOTs, and readout of the Néel order vector in AFM CuMnAs layers, via anisotropic magnetoresistance [11] and Voigt effect [23], was an important step toward next-generation AFM-based solid-state memory chips. It was also shown that the fourfold degeneracy of Dirac points in this material can be lifted by manipulating nonsymmorphic symmetries. This is achieved by electrically altering the direction of the Néel vector through SOTs. Therefore a topological metal-insulator transition can happen in this system electrically [6].
@@
AC conductivity is a crucial parameter in condensed matter physics because it provides information on the electronic and magnetic behavior of materials in response to electromagnetic radiation, particularly in the optical frequency range [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. AC conductivity is related to various material properties such as dielectric functions, absorption spectra, charge susceptibilities, plasmon dispersion, and its linewidth, as well as magneto-optical effects [36, 37, 38, 24]. Very recently, terahertz time-domain spectroscopy was used to measure the anisotropic conductivity of CuMnAs thin films [39]. Although there are a few studies in optical properties of 2D Dirac materials in the presence of an external magnetic field with broken $\mathcal{PT}$ symmetry [40, 41], to the best of our knowledge, a comprehensive investigation of the AC conductivity of 2D AFM CuMnAs layers that preserve $\mathcal{PT}$ symmetry has not been conducted so far. This article aims to bridge this gap by examining the AC conductivity, magnetic dichroism, and dielectric function of the 2D AFM CuMnAs layer. We specifically explore variations in these properties by tuning the orientation of the Néel vector by employing a time-dependent linear response Kubo formalism.
@@ II Model
In the presence of AFM order, it is possible to break the fourfold degeneracy of Dirac points and open a gap and have a semimetal-insulator transition. The condition for band crossing, $E_{\bm{k}}^{+}=E_{\bm{k}}^{-}$, requires $\varepsilon_{\bm{k}}=0$ and thus $\gamma_{1}=\delta_{x}=\delta_{y}=\delta_{z}=0$. To preserve the fourfold degeneracy at the Dirac points, where $\gamma_{1}=0$, in the presence of AFM order, the out-of-plane component of the Néel vector must be zero $n_{z}=0$. In addition, the system must satisfy either of the following two criteria: $(i)$ at $k_{x}=\pm\pi/a$, we should have $\sin(ak_{y})=J_{\rm{n}}/\lambda\leq 1$, that requires $n_{x}=1$ and $n_{y}=0$; or $(ii)$ at $k_{y}=\pm\pi/a$, we should have $\sin(ak_{x})=J_{\rm{n}}/\lambda\leq 1$, that requires $n_{x}=0$ and $n_{y}=1$. Therefore, the only directions of the Néel vector in which the fourfold degeneracy is preserved at two out of three Dirac points, are inplane directions along the $x$ and $y$ directions if the exchange coupling is smaller than the SOC parameter, $J_{\rm{n}}/\lambda\leq 1$. It is evident that by changing the direction of the Néel vector, one can turn on and off the Dirac gap in this system. In addition, the positions of the Dirac cones are shifted by changing the direction of the Néel vector, see Fig. 2. When the Néel vector has either an out-of-plane or inplane component between the $x$ and $y$ directions, all Dirac points become gapped, see electronic dispersion with the green and black curves in bottom panel of Fig. 2. Conversely, if the Néel vector lies inplane along the $x$ and $y$ directions, only some of Dirac points become gapped in 1st BZ, as shown by the blue and red curves in Fig. 2. With the material parameters used in Fig. 2, the system becomes an insulator when the Néel vector is aligned along the $z$ direction.
@@ III AC conductivity
Figure 3: The real (top) and imaginary (bottom) parts of the longitudinal conductivity along the $x$ direction in the absence (gray color line) and presence of the AFM order as a function of the frequency. The conductivity is changed by varying the Néel vector direction. There is no DC conductivity when the Néel vector is out-of-plane since the Fermi energy lies inside the electronic band gap and the system becomes an insulator. $\hbar\eta/E_{F}=0.002$
The conductivity tensor satisfies the Onsager-Büttiker relation $\sigma_{\alpha\beta}(\omega;\bm{n})=\sigma_{\beta\alpha}(\omega;-\bm{n})$. If the system exhibits $\mathcal{PT}$ symmetry, the conductivity tensor contains only symmetric components, and there is no Hall effect. However, a transverse response and an anisotropic longitudinal response are still possible, depending on the relative direction of the applied electric field and the Néel vector in the presence of SOC. In the absence of AFM order or in the presence of an out-of-plane Néel vector, the off-diagonal elements of the conductivity tensor vanish and the longitudinal conductivity is isotropic. In general, if the Néel vector resides in the $xz$ and $yz$ planes, the transverse conductivity is zero and if the Néel vector has an inplane component, the longitudinal conductivity is anisotropic.
To better understand direct optical transitions around three Dirac points in the absence of the AFM order, we first linearize the dispersion around these three points. Around the $\bm{M}=(\pi,\pi)$ symmetry point, the low-energy gapless Dirac-like dispersion reads,
@@
Figure 4: The real (top) and imaginary (bottom) parts of the transverse conductivity in the presence of the AFM order as a function of the frequency. The sign and amplitude of the transverse conductivity are changed by varying the Néel vector direction. The transverse conductivity is zero when the Néel vector direction is in the $xz$ or $yz$ planes. We set $\hbar\eta/E_{F}=0.002$.
@@ IV Magnetic dichroism
Magnetic circular and linear dichroism are optical techniques for measuring the electronic band structure and the magnetic orientation of the system [24]. These two effects are related to different components of the AC conductivity tensor. Circular dichroism is the differential absorption of left and right circularly polarized light in a magnetically ordered material. This effect is finite in chiral systems in which antisymmetric part of the off-diagonal elements of the AC conductivity tensor is finite [41, 43, 44, 45, 46]. In the Dirac AFM system, considered in this study, because of the $\mathcal{PT}$ symmetry, there is no antisymmetric (anomalous Hall) component and thus no circular dichroism. Therefore, both magneto-optical Faraday and Kerr effects are absent in these materials. This fact makes optical readout of the order parameter direction seem a challenging task.
Figure 5: The magnetic linear dichroism (LD) as a function of the frequency, Eq. (9). The amplitude and the sign of the LD signal is sensitive to the direction of the Néel vector and frequency.
On the other hand, the magnetic linear dichroism is the difference between the absorption of light polarized parallel and that polarized perpendicular to an orientation axis. This effect is related to the anisotropy in the longitudinal components of the AC conductivity tensor. While the Faraday and Kerr effects are first-order in relation to the local magnetic moment, linear dichroism can arise from the Voigt-Cotton–Mouton effect [47], which is proportional to the square of the local magnetic moment. Additionally, it may stem from electronic anisotropy, associated with symmetry breaking due to changes in the direction of the Néel vector [48, 49]. Hence, the linear dichroism is a powerful tool for directly exploring the symmetry breaking in the system. The linear dichroism is defined as,
@@ V Dielectric function and reflective index
where $\epsilon_{0}$ is the static permittivity of the substrate and $\sigma_{\pm}=\big{[}(\sigma_{xx}+\sigma_{yy})\pm i(\sigma_{xy}-\sigma_{yx})\big{]}/2$ is the AC conductivity for the left (-) and right (+) circularly polarized light. Since there is no antisymmetric contribution in the transverse conductivity tensor, the dielectric function is the same for both polarizations $\epsilon_{+}=\epsilon_{-}$, and thus $n_{+}=n_{-}$. Therefore there is no circular birefringence in the system under study. However, since in general $n_{xx}\neq n_{yy}$ in the system, there might be a finite linear birefringence [47], which can be tuned by the Néel vector orientation.
Figure 6, represents the real (top panel) and imaginary (bottom panel) parts of the dielectric function and show the tuning of the Néel vector changes the dielectric response and thus optical properties of the system.
@@ Acknowledgements
R.S. and S.E. would like to acknowledge financial support from the research council of University of Tehran.
A.Q. acknowledges financial support from the Research Council of Norway through its Centers of Excellence funding scheme, project number 262633, ”QuSpin”.

### v2

@@ I Introduction
On-demand manipulation and readout of the staggered Néel order parameter in antiferromagnetic (AFM) systems are challenging problems in the emerging field of AFM-based spintronics [1, 2]. Both theoretically and experimentally, spin-orbit and spin-transfer torques have been demonstrated to be finite in AFM systems under certain symmetry conditions [3, 4, 5, 6, 7, 8, 9, 10, 11]. It was shown that in AFM systems with preserved combined $\mathcal{PT}$ symmetry, while the time-reversal $\mathcal{T}$ symmetry and the inversion $\mathcal{P}$ symmetry are each broken, staggered spin-orbit torques acting on the magnetic sublattices may reorient the AFM Néel vector [11, 6, 5, 10, 3, 12].
@@
Two-dimensional (2D) Dirac materials are a recently discovered class of 2D materials that are not gapped by spin-orbit couplings (SOCs) [13, 14, 15], and thus exhibit behavior distinct from graphene [16, 17], surface states of topological insulators [18], and 3D Dirac and Weyl semimetals [19, 20]. They have more than one Dirac point in their first Brillouin zone (BZ). The band crossing at the Dirac points of these materials is protected by nonsymmorphic lattice symmetry [13]. The 2D CuMnAs is a prototype of Dirac materials with a long-range AFM order [21, 22, 23]. The recent discovery of electrical switching, via staggered spin-orbit torque, and readout of the Néel order vector in CuMnAs AFM layers, via anisotropic magnetoresistance [11, 24] and Voigt effect [25], was an important step towards next generation AFM-based solid-state memory chips. It was also shown that the fourfold degeneracy of Dirac points in this material can be lifted by manipulating nonsymmorphic symmetries. This is achieved by electrically altering the direction of the Néel vector through spin-orbit torques. Therefore a topological metal-insulator transition can happen in this system electrically [6].
@@
AC conductivity is a crucial parameter in condensed matter physics because it provides information on the electronic and magnetic behavior of materials in response to electromagnetic radiation, particularly in the optical frequency range [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. AC conductivity is related to various material properties such as dielectric functions, absorption spectra, charge susceptibilities, plasmon dispersion, and its linewidth, as well as magneto-optical effects [38, 39, 40, 26]. Very recently, terahertz time-domain spectroscopy was used to measure the anisotropic conductivity of CuMnAs thin films [41]. Although there are a few studies on the optical properties of 2D Dirac materials in the presence of an external magnetic field with broken $\mathcal{PT}$ symmetry [42, 43], to the best of our knowledge, no comprehensive investigation of the AC conductivity of 2D AFM CuMnAs layers that preserve $\mathcal{PT}$ symmetry [24] has been conducted so far [12]. This article aims to bridge this gap by examining the AC conductivity, magnetic dichroism, and dielectric function of the 2D AFM CuMnAs layer. We specifically explore variations in these properties by tuning the orientation of the Néel vector by employing a time-dependent linear response Kubo formalism.
@@ II Model
In the presence of AFM order, it is possible to break the fourfold degeneracy of Dirac points and open a gap and have a semimetal-insulator transition. The condition for band crossing, $E_{\bm{k}}^{+}=E_{\bm{k}}^{-}$, requires $\varepsilon_{\bm{k}}=0$ and thus $\gamma_{1}=\delta_{x}=\delta_{y}=\delta_{z}=0$. To preserve the fourfold degeneracy at the Dirac points, where $\gamma_{1}=0$, in the presence of AFM order, the out-of-plane component of the Néel vector must be zero $n_{z}=0$. In addition, the system must satisfy either of the following two criteria: $(i)$ at $k_{x}=\pm\pi/a$, we should have $\sin(ak_{y})=J_{\rm{n}}/\lambda\leq 1$, that requires $n_{x}=1$ and $n_{y}=0$; or $(ii)$ at $k_{y}=\pm\pi/a$, we should have $\sin(ak_{x})=J_{\rm{n}}/\lambda\leq 1$, that requires $n_{x}=0$ and $n_{y}=1$. Therefore, the only directions of the Néel vector in which the fourfold degeneracy is preserved at two out of three Dirac points, are inplane directions along the $x$ and $y$ directions if the exchange coupling is smaller than the SOC parameter, $J_{\rm{n}}/\lambda\leq 1$. It is evident that by changing the direction of the Néel vector, one can turn on and off the Dirac gap in this system. In addition, the positions of the Dirac cones are shifted by changing the direction of the Néel vector, see Fig. 2. When the Néel vector has either an out-of-plane or inplane component between the $x$ and $y$ directions, all Dirac points become gapped; see electronic dispersion with the green and black curves in the bottom panel of Fig. 2. Conversely, if the Néel vector lies inplane along the $x$ or $y$ directions, only some of Dirac points become gapped in 1st BZ, as shown by the blue and red curves in Fig. 2. With the material parameters used in Fig. 2, the system becomes an insulator when the Néel vector is aligned along the $z$ direction.
@@ III AC conductivity
Figure 3: The real (top) and imaginary (bottom) parts of the longitudinal conductivity along the $x$ direction in the absence (gray color line) and presence of the AFM order as a function of the frequency. The conductivity is changed by varying the Néel vector direction. There is no DC conductivity when the Néel vector is out-of-plane since the Fermi energy lies inside the electronic band gap and the system becomes an insulator. $\hbar\eta=0.002E_{F}$
The conductivity tensor satisfies the Onsager-Büttiker relation $\sigma_{\alpha\beta}(\omega;\bm{n})=\sigma_{\beta\alpha}(\omega;-\bm{n})$. If the system exhibits $\mathcal{PT}$ symmetry, the conductivity tensor contains only symmetric components, and there is no Hall effect. However, a transverse response and an anisotropic longitudinal response are still possible, depending on the relative direction of the applied electric field and the Néel vector in the presence of SOC. In the absence of AFM order or in the presence of an out-of-plane Néel vector, the off-diagonal elements of the conductivity tensor vanish and the longitudinal conductivity is isotropic. In general, if the Néel vector resides in the $xz$ or $yz$ planes, the transverse conductivity is zero. When the Néel vector has an inplane component, the longitudinal conductivity is anisotropic.
To better understand direct optical transitions around three Dirac points in the absence of the AFM order, we first linearize the dispersion around these three points. Around the $\bm{M}=(\pi,\pi)$ symmetry point, the low-energy gapless Dirac-like dispersion is isotropics and reads,
@@
Figure 4: The real (top) and imaginary (bottom) parts of the transverse conductivity in the presence of the AFM order as a function of the frequency. The sign and amplitude of the transverse conductivity are changed by varying the Néel vector direction. The transverse conductivity is zero when the Néel vector direction is in the $xz$ or $yz$ planes. We set $\hbar\eta=0.002E_{F}$.
@@ IV Magnetic dichroism
Magnetic circular and linear dichroism are optical techniques for measuring the electronic band structure and the magnetic orientation of the system [26]. These two effects are related to different components of the AC conductivity tensor. Circular dichroism is the differential absorption of left and right circularly polarized light in a magnetically ordered material. This effect is finite in chiral systems in which antisymmetric part of the off-diagonal elements of the AC conductivity tensor is finite [43, 45, 46, 47, 48]. In the Dirac AFM system, considered in this study, because of the $\mathcal{PT}$ symmetry, there is no antisymmetric (anomalous Hall) component [24] and thus no circular dichroism. Therefore, both magneto-optical Faraday and Kerr effects are absent in this system. This fact makes the optical readout of the staggered order parameter direction challenging.
Figure 5: The magnetic linear dichroism (LD) as a function of the frequency, Eq. (9). The amplitude and sign of the LD signal are sensitive to the direction of the Néel vector and frequency.
On the other hand, the magnetic linear dichroism is the difference between the absorption of light polarized parallel and that polarized perpendicular to an orientation axis. This effect is related to the anisotropy in the longitudinal components of the AC conductivity tensor. Although the Faraday and Kerr effects are first order in relation to the local magnetic moment, linear dichroism may arise from the Voigt-Cotton–Mouton effect [49], which is proportional to the square of the local magnetic moment. Additionally, it may stem from electronic anisotropy, associated with symmetry breaking due to changes in the direction of the Néel vector [50, 51]. Hence, linear dichroism is a powerful tool for directly exploring symmetry breaking in the system. The linear dichroism is defined as,
@@ V Dielectric function and reflective index
where $\epsilon_{0}$ is the static permittivity of the substrate and $\sigma_{\pm}=\big{[}(\sigma_{xx}+\sigma_{yy})\pm i(\sigma_{xy}-\sigma_{yx})\big{]}/2$ is the AC conductivity for the left (-) and right (+) circularly polarized light. Since there is no antisymmetric contribution in the transverse conductivity tensor, the dielectric function is the same for both polarizations $\epsilon_{+}=\epsilon_{-}$, and thus $n_{+}=n_{-}$. Therefore there is no circular birefringence in the system under study. However, since in general $n_{xx}\neq n_{yy}$ in this system, there might be a finite linear birefringence [49], which can be tuned by the Néel vector orientation.
Figure 6, represents the real (top panel) and imaginary (bottom panel) parts of the dielectric function and shows that the tuning of the Néel vector changes the dielectric response and thus the optical properties of the system.
@@ Acknowledgements
R.S. and S.E. would like to acknowledge financial support from the research council of the University of Tehran. A.Q. acknowledges financial support from the Research Council of Norway through its Centers of Excellence funding scheme, project number 262633, “QuSpin”.